The area of a rectangle is 21sqft. What are the dimensions is the length is 5 less four times the width.
Found 2 solutions by nyc_function, user_dude2008:
Answer by nyc_function(2741)

Here is your equation and you can then solve the problem.
21 = (4x - 5)(x)

Answer by user_dude2008(1862)

21=LW
L=4W-5

21=(4W-5)W

4W^2-5W-21=0

(W-3)(4W+7)=0

W-3=0 or 4W+7=0

W=3 or W=-7/4 Ignore

Answer: Width = 3, length = 7